# Two-Phase Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v2

@@ Abstract
Reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is emerging as a key technology to support ultra-high data rates in future sixth-generation networks. However, the acquisition of accurate channel state information (CSI) in such systems is challenging due to the passive nature of RIS and the hybrid beamforming architecture typically employed in THz systems. To address these challenges, we propose a novel low-complexity two-phase channel estimation scheme for RIS-assisted THz systems with beam split effect. In the proposed scheme, we first estimate the full CSI over a small subset of subcarriers, then extract angular information at both the base station and RIS. Subsequently, we recover the full CSI across remaining subcarriers by determining the corresponding spatial directions and angle-excluded coefficients. Theoretical analysis and simulation results demonstrate that the proposed method achieves superior performance in terms of normalized mean-square error while significantly reducing computational complexity compared to existing algorithms.
@@ I Introduction
Envisioned as a cornerstone technology to meet the requirements of future 6G networks, terahertz (THz) communication holds immense potential for a myriad of applications within the field of the Internet of Things, ranging from intelligent industrial manufacturing to digital twinning technologies and immersive experiences like holographic telepresence and the metaverse [1, 2, 3]. However, THz communication systems encounter various challenges, including high propagation loss, limited coverage, and susceptibility to atmospheric absorption and scattering [4, 5]. One notable drawback is the severe signal attenuation experienced by transmitted signals, which often necessitate a line-of-sight (LoS) link to maintain communication quality [6]. To address these challenges and unlock the full potential of THz communication, reconfigurable intelligent surface (RIS) has emerged to provide LoS paths by intelligently manipulating the propagation environment [7]. RIS can dynamically adjust reflection properties of a large number of passive reflecting elements to optimize signal transmission and reception [8]. By strategically controlling the phase shifts of these reflecting elements, RIS can effectively enhance the signal strength, mitigate interference, and improve the overall performance of THz communication systems. However, due to the passive nature of RIS, which lacks signal processing capabilities, the performance enhancement enabled by RIS relies significantly on accurate channel state information (CSI), the acquisition of which can be challenging. Therefore, investigating an efficient channel estimation scheme for RIS-assisted THz communication systems is of great importance.
I-A Prior Work
@@
Over the past few years, numerous research works have been dedicated to exploring channel estimation in RIS-assisted high frequency band systems [9, 10, 11, 12, 13, 14]. Specifically, a sparse representation of RIS-assisted millimeter wave (mmWave) channel has been established, on which compressed sensing based techniques can be implemented for CSI acquisition [9]. In [10], a low-overhead channel estimation method has been proposed based on the double-structure characteristic of RIS-assisted angular cascaded channel. To reduce computational complexity, [11] proposed a non-iterative two-stage framework based on direction-of-arrival (DoA) estimation. The authors in [12] proposed a two-phase channel estimation strategy to reduce pilot overhead. [13] investigated cooperative-beam-training based channel estimation procedure. In [14], a sparse Bayesian learning based channel estimation approach has been developed to compensate the severe off-grid issue.
@@
The above solutions have effectively addressed the channel estimation problem of narrowband systems with utilization of the angular domain sparsity. However, THz communication typically adopts a hybrid beamforming architecture to reduce hardware costs[15]. When considering this architecture, the beam split effect becomes inevitable. This effect occurs due to the wider bandwidth of THz signals and the enormous number of RIS elements, which causes path components to split into distinct spatial directions at different subcarrier frequencies [16]. This phenomenon leads to a considerable degradation in achievable data rates, which counteracts the benefits of RIS-assisted THz systems.
@@
To address this issue, beam split effect in RIS-assisted THz systems has been investigated in [17, 18, 19]. In particular, [17] introduced a beam-split-aware orthogonal matching pursuit (BSA-OMP) approach to automatically mitigate beam split. In [18], a low-complexity beam squint mitigation scheme is proposed based on generalized-OMP estimator. The authors in [19] proposed a robust regularized sensing-beam-split-OMP-based scheme to estimate beam split affected RIS-assisted cascaded channel. These works primarily focus on mitigating beam split effects within the context of cascaded channels using well-designed dictionaries covering the entire beamspace. However, an alternative approach is to focus on the physical directions of the channel components to reduce computational complexity [20, 21, 22]. Specifically, the authors in [20] regarded channel estimation problem as a block-sparse recovery problem. [21] proposed a beam split pattern detection based wideband channel estimation scheme where the physical DoAs are prioritized for determination. In [22], a three-stage wideband channel estimation scheme is proposed where angle-of-arrivals (AoAs) and angle-of-departure (AoDs) are determined and refined to reconstruct the wideband channel with beam split effect. Although these algorithms achieve accurate channel estimation, they require execution at each SC or simultaneous processing of all SCs, which may introduce additional unaffordable computations. Moreover, the passive nature of RIS complicates the precise acquisition of angles at the RIS, which often results in the availability of only coupled angles. Therefore, this motivates the development of a low-complexity channel estimation scheme in RIS-assisted wideband THz systems with beam split effect.
@@ I-B Main Contributions
To address the issue discussed above, we presented some preliminary results on estimating the RIS-assisted cascaded channel [23]. In this paper, we significantly expand on the previous work and propose a two-phase efficient channel estimation scheme for RIS-assisted wideband THz systems with beam split. Our main contributions are summarized as follows:
  • We propose a novel low-complexity channel estimation protocol for general RIS-assisted wideband THz communication scenarios with beam split effect, as depicted in Fig. 1. The channel estimation scheme is divided into two phases: full CSI estimation for two SCs in Phase I and remaining SCs estimation in Phase II. By employing information from only two SCs, this protocol can significantly reduce the computational complexity while maintaining complete and accurate channel estimation.
  • In Phase I, we propose a two-step method to acquire full CSI for the selected two SCs. Specifically, in Step 1, we propose a simplified cyclic beam split (CBS) dictionary to reduce the size from the scale of $Q^{2}$ to a linear scale of $Q$. Using this dictionary, we reformulate the problem as a generalized sparse recovery problem. Moreover, we discuss the selection of different sparse recovery algorithms. In Step 2, we first design a heuristic DoD estimation method at BS according to the estimated cascaded channel obtained in Step 1. We then reformulate the coupled angles at RIS as a typical DoA estimation problem and design a double sensing-multiple signal classification (DS-MUSIC) algorithm by integrating the angular information from the two selected SCs. We also provide insights into the selection of two SCs. In Phase II, given angular information from Phase I, we employ a simple least squares (LS) algorithm to estimate the channels for the remaining SCs.
  • Through the analysis of algorithm complexity, our proposed protocol significantly reduces the computational complexity compared to the existing sparse recovery solutions. Additionally, numerical results demonstrate that the NMSE performance of our proposed channel estimation scheme closely approaches that of oracle-LS method. By focusing on a two-step estimation process for two SCs and applying a simplified LS algorithm for the remaining SCs, the proposed method can achieve both reduction of computational complexity and acquisition of accurate CSI.
Figure 1: Two-phase channel estimation protocol and frame structure.
@@
The remainder of this paper is organized as follows. In Section II, we first describe the RIS-assisted wideband THz communication setup and the beam split effect. In Section III, we describe our proposed channel estimation protocol. Section V provides an analysis of the computational complexity of our proposed scheme, with a comparison to existing methods. Finally, we conclude the paper in Section VI.
@@
Notations: Boldface lower-case and capital letters represent column vectors and matrices, respectively. $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\mathsf{T}}$, $\left(\cdot\right)^{\mathsf{H}}$ and $\left(\cdot\right)^{\dagger}$ denote conjugate, transpose, transpose-conjugate and pseudo-inverse operation, respectively. $\mathbb{E}$ and $\mathrm{Var}$ represent statistical expectation and variance, respectively. $\mathbb{N}$, $\mathbb{C}$ and $\mathbb{R}$ denote the set of integers, complex numbers and real numbers, respectively. $\mathbf{I}_{N}$ is $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace. $\otimes$ denote Kronecker product, * and $\bullet$ denote Khatri-Rao product and transposed Khatri-Rao product, respectively.
@@ II-A Signal and Channel Model
Figure 2: RIS-assisted THz communication system with beam split effect.
@@
We consider a RIS-assisted wideband THz MIMO system employing hybrid beamforming architecture that integrates both analog and baseband processing, as depicted in Fig. 2. Assuming that the BS equipped with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}$ (where $N_{\mathrm{RF}}<N_{\mathrm{T}}$) RF chains is deployed to serve $K$ single-antenna UEs. The RIS employs a uniform linear array (ULA) comprised of $N_{\mathrm{R}}$ reflecting elements, which is manipulated by a smart controller operated by the BS. Besides, the OFDM system is considered with $M$ subcarriers, central frequency $f_{c}$ and bandwidth $B$.
@@
In this paper, we assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [12]. For each coherence block, channel estimation is firstly performed during $T$ subframes, and downlink data transmission is then carried out during the remaining $D-T$ subframes. Moreover, each subframe of the downlink channel estimation period is divided into $P$ time slots, with $P$ baseband beamformers operating alternatively across different time slots. Additionally, we assume that the direct BS-UE links are blocked by obstacles [20]. Therefore, we mainly focus on the downlink channel estimation of the cascaded BS-RIS-UE channel.
@@
As a cost-performance trade-off, hybrid beamforming architecture is typically employed in THz MIMO systems[15]. In our considered system, we assume that $P$ frequency-dependent baseband beamformers $\mathbf{W}_{t}^{\mathrm{BB}}[m]=\left[\mathbf{w}^{\mathrm{BB}}_{t,1}[m],\ldots,\mathbf{w}^{\mathrm{BB}}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{RF}}\times P}$ are employed to process $P$ orthogonal pilot signals $\mathbf{S}_{t}\triangleq\mathrm{Dia{\tiny}g}\left\{s_{t,1},\cdots,s_{t,P}\right\}\in\mathbb{C}^{P\times P}$ per time slot. Meanwhile, a frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$ is employed to achieve gain improvement in the direction of RIS with each entry is subject to the unit-modulus constraint, i.e., $|[\mathbf{W}^{\mathrm{RF}}]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ for $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$. We assume that the analog beamformer remains unchanged over multiple channel coherence blocks due to the relatively static nature of the RIS physical location compared to the dynamic channel coefficients. Therefore, we define $\mathbf{W}_{t}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{W}_{t}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$ as the hybrid beamformer which consists of $P$ beamformer vectors, i.e., $\mathbf{W}_{t}[m]=\left[\mathbf{w}_{t,1}[m],...,\mathbf{w}_{t,P}[m]\right]$. Let $\mathbf{G}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ and $\mathbf{h}_{k}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}}$ denote the channel from BS to RIS and RIS to $k$-th UE at the $m$-th SC, respectively. At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by $k$-th UE at the $m$-th SC can be expressed as
$\begin{split}y_{k,t,p}[m]\!&=\!\mathbf{h}_{k}\left[m\right]\mathrm{Diag}\left(\boldsymbol{\theta}_{t}\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m]\\ &=\!\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m],\end{split}$ (1)
@@
where $n_{k,t,p}[m]\sim\mathcal{CN}\left(0,2\varepsilon\right)$ is the complex additive white Gaussian noise (AWGN). Besides, $\boldsymbol{\theta}_{t}=\left[e^{j\theta_{t,1}},\cdots,e^{j\theta_{t,N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS at subframe $t$ with a constant reflection amplitude constraint [24].
@@
Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t,p}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
@@
$\mathbf{G}[m]=\sum_{l=1}^{L}{\alpha_{l}\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{l})\mathbf{a}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\varphi}_{l}\right)e^{-j2\pi\tau_{l}f_{m}}},$ (4)
where $L$ represents the number of propagation paths between BS and RIS, $\alpha_{l}$, $\tau_{l}$, $\tilde{\psi}_{l}$ and $\tilde{\varphi}_{l}$ represent the complex gain, the time delay, the physical angle-of-departure (AoD) at BS and angle-of-arrival (AoA) at RIS for the $l$-th path, respectively. Besides, $f_{m}$ denotes the frequency of the $m$-th SC, which is given by
@@
$\mathbf{a}_{\mathrm{T},m}(\tilde{u})\triangleq\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j2\pi\sin\left(\tilde{u}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
$\mathbf{a}_{\mathrm{R},m}(\tilde{v})\triangleq\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j2\pi\sin\left(\tilde{v}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th suncarrier. $d_{\mathrm{BS}}$ and $d^{\mathrm{RIS}}$ denote the antenna spacing of BS and RIS, respectively. Besides, $\mathbf{n}_{t}=[0,1,\dots,N_{\mathrm{T}}-1]$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]$, respectively. By defining the angle vectors $\tilde{\boldsymbol{\psi}}=[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be transformed into the virtual angular representation (VAD) form as

### v3

@@ Abstract
Reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is emerging as a key technology to support the ultra-high data rates in future sixth-generation networks. However, the acquisition of accurate channel state information (CSI) in such systems is challenging due to the passive nature of RIS and the hybrid beamforming architecture typically employed in THz systems. To address these challenges, we propose a novel low-complexity two-phase channel estimation scheme for RIS-assisted THz systems with beam split effect. In the proposed scheme, we first estimate the full CSI over a small subset of subcarriers (SCs), then extract angular information at both the base station and RIS. Subsequently, we recover the full CSI across remaining SCs by determining the corresponding spatial directions and angle-excluded coefficients. Theoretical analysis and simulation results demonstrate that the proposed method achieves superior performance in terms of normalized mean-square error while significantly reducing computational complexity compared to existing algorithms.
@@ I Introduction
Envisioned as a cornerstone technology to meet the requirements of future 6G networks, terahertz (THz) communication holds immense potential for a myriad of applications within the field of the Internet of Things, ranging from intelligent industrial manufacturing to digital twinning technologies and immersive experiences like holographic telepresence and the metaverse [1, 2, 3]. However, THz communication systems encounter various challenges, including high propagation loss, limited coverage, and susceptibility to atmospheric absorption and scattering [4, 5]. One notable drawback is the severe signal attenuation experienced by transmitted signals, which often necessitate a line-of-sight (LoS) link to maintain communication quality [6]. To address these challenges and unlock the full potential of THz communication, reconfigurable intelligent surface (RIS) has emerged to provide virtual LoS path by intelligently manipulating the propagation environment [7]. RIS can dynamically adjust reflection properties of a large number of passive reflecting elements to optimize signal transmission and reception [8]. By strategically controlling the phase shifts of these reflecting elements, RIS can effectively enhance the signal strength, mitigate interference, and improve the overall performance of THz communication systems. Despite these benefits, the RIS elements remain passive and lack inherent signal processing capabilities. Accordingly, the performance enhancement enabled by RIS relies significantly on the accurate channel state information (CSI), which is challenging to acquire. Therefore, investigating an efficient channel estimation scheme for RIS-assisted THz communication systems is of great importance.
I-A Related Works
@@
Over the past few years, numerous research works have been dedicated to exploring channel estimation in RIS-assisted high frequency systems [9, 10, 11, 12, 13, 14]. Specifically, a sparse representation of RIS-assisted millimeter wave (mmWave) channel has been established, enabling the application of compressed sensing based techniques for CSI acquisition [9]. In [10], a low-overhead channel estimation method has been proposed based on the double-structure characteristic of RIS-assisted angular cascaded channel. To reduce computational complexity, [11] proposes a non-iterative two-stage framework based on direction-of-arrival (DoA) estimation. The authors in [12] propose a two-phase channel estimation strategy to reduce pilot overhead. [13] investigates cooperative-beam-training based channel estimation procedure. In [14], a sparse Bayesian learning based channel estimation approach has been developed to compensate the severe off-grid issue.
@@
The above solutions have effectively addressed the channel estimation problem of narrowband systems with utilization of the angular domain sparsity. However, mmWave/THz communications typically adopt a hybrid beamforming architecture to reduce hardware costs[15], making beam squint/beam split effect inevitable. These effects occur due to the wider bandwidth of mmWave/THz signals and the enormous number of RIS elements, which causes path components to split into distinct spatial directions at different subcarrier (SC) frequencies [16]. Such phenomenon significantly degrades achievable data rates, thereby counteracting the benefits of RIS-assisted THz systems.
@@
To address this issue, beam squint/beam split effect in RIS-assisted mmWave/THz systems has been investigated in [17, 18, 19, 20]. In particular, [17] introduces a beam-split-aware orthogonal matching pursuit (BSA-OMP) approach to automatically mitigate beam split effect. In [18], a low-complexity beam squint mitigation scheme is proposed based on generalized-OMP estimator. The authors in [19] propose a robust regularized sensing-beam-split-OMP-based scheme to estimate beam split affected RIS-assisted cascaded channel. The authors in [20] regard the wideband channel estimation problem as a block-sparse recovery problem by concatenating signals at different SCs. These works primarily focus on mitigating beam split effect within the context of cascaded channels using well-designed dictionaries covering the entire beamspace. Although they can achieve accurate channel estimation, it requires execution at each SC or simultaneous processing of all SCs, which may introduce additional unaffordable computations.
An alternative approach is to focus on the physical directions of the channel components to reduce computational complexity [21, 22, 23, 24, 25]. In mmWave band, angle-based channel estimation methods are used to simplify the wideband channel estimation process under the beam squint effect [21, 22]. The authors in [21] first estimate angle-of-arrivals (AoAs) at BS and then employed BSA-based algorithms to acquire full CSI. [22] presents a twin-stage OMP method to estimate the path angles of the RIS-assisted cascaded two-hop channel. [23] introduces a Newtonized OMP (NOMP) based algorithm to detect the channel parameters. These approaches perform angle estimation by leveraging the principle that the frequency-dependent spatial angles change while frequency-independent physical angles remain constant across different SCs. This principle can also be applied in THz systems, however, due to the significantly larger beam spacing in THz band compared to mmWave systems, the frequency-dependent spatial angles exhibit greater deviations. Moreover, the solutions discussed above fail to leverage the wider bandwidth available in THz band to better differentiate between true physical and false spatial angles. Consequently, the angle-based methods used in mmWave systems may become neither efficient nor effective when applied to the THz band.
To this end, several low-complexity estimation schemes have been developed specifically for wideband THz systems [24, 25]. For example, [24] proposes a beam-split-pattern detection-based estimation method that prioritizes physical DoA determination. In [25], a three-stage wideband channel estimation approach is presented, where AoAs and angles-of-departure (AoDs) are successively refined to reconstruct the channel under beam split effect. These methods achieve a favorable trade-off between complexity and performance in traditional architectures. However, these methods are not readily applicable to RIS-assisted THz systems due to the lack of mechanisms for angular estimation at RIS. Moreover, the passive nature of RIS further complicates the accurate acquisition of angular information, making it difficult to resolve the coupled AoAs/AoDs. Therefore, this motivates the development of a low-complexity channel estimation scheme in RIS-assisted wideband THz systems with beam split effect.
Figure 1: Illustration of the two-phase channel estimation protocol and corresponding frame structure.
@@ I-B Main Contributions
To address the issues discussed above, we presented some preliminary results on estimating the RIS-assisted cascaded channel [26]. In this paper, we significantly expand on the previous work and propose a two-phase efficient channel estimation scheme for RIS-assisted wideband THz systems with beam split. Our main contributions are summarized as follows:
  • We propose a novel low-complexity channel estimation scheme tailored for RIS-assisted wideband THz communication systems under the beam split effect. Unlike conventional methods that rely on full-band sparse recovery across all subcarriers, the proposed scheme adopts a two-phase protocol to achieve accurate CSI estimation with substantially reduced computations. Specifically, as illustrated in Fig. 1, Phase I estimates the full CSI at two carefully selected SCs, while Phase II efficiently reconstructs the remaining CSI using only a simple least squares (LS) algorithm, thereby avoiding the need of computationally expensive recovery procedures across the full bandwidth.
  • The proposed low-complexity design is primarily enabled by the two-step approach implemented in Phase I. In Step 1, we recover the full CSI at two selected SCs utilizing a simplified cyclic beam split (CBS) dictionary, which is compressed from the original size of $Q^{2}$ to a linear scale of $12Q-$ 1, thereby significantly reducing the complexity of sparse recovery. In Step 2, we estimate the physical directions at both the BS and RIS based on the recovered CSI. Specifically, we propose an energy maximum (EnM) algorithm for directions-of-departure (DoDs) estimation at BS and a double sensing-MUSIC (DS-MUSIC) algorithm to resolve the coupled angles at RIS. Instead of mitigating the severe beam misalignment introduced by beam split effect, the proposed algorithm leverages the resulting spatial separation of RIS angles within the extended angular range, which facilitates accurate estimation of physical directions.
  • We provide a comprehensive complexity analysis, which demonstrates that our proposed protocol significantly reduces computational complexity compared to existing sparse recovery solutions. Furthermore, we conduct extensive numerical evaluations under both mmWave and THz system settings. The results demonstrate that: (i) the proposed two-phase scheme efficiently exploits the beam split properties to enhance estimation accuracy compared to mmWave solutions; and (ii) it achieves superior NMSE performance over conventional baselines, approaching that of the oracle-LS benchmark. These findings highlight the potential of our method for efficient channel estimation in RIS-assisted THz communication systems.
@@
The remainder of this paper is organized as follows. Section II presents the system model for RIS-assisted wideband THz communication with beam split effect along with the associated channel estimation protocol. Section III, IV and V describe the proposed two-phase channel estimation framework in detail. Section VI analyzes the computational complexity, and Section VII provides simulation results and performance comparisons with existing methods. Finally, Section VIII concludes the paper.
@@
Notations: Boldface lower-case and capital letters represent the column vectors and matrices, respectively. $\left(\cdot\right)^{*}$, $(\cdot)^{\mathsf{T}}$, $\left(\cdot\right)^{\mathsf{H}}$ and $\left(\cdot\right)^{\dagger}$ denote the conjugate, transpose, transpose-conjugate and pseudo-inverse operation, respectively. $\mathbb{E}$ and $\mathrm{Var}$ represent the statistical expectation and variance, respectively. $\mathbb{N}$, $\mathbb{C}$ and $\mathbb{R}$ denote the set of integers, complex numbers and real numbers, respectively. $\mathbf{I}_{N}$ is the $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace. $\otimes$, $\odot$ and $\bullet$ denote the Kronecker product, Khatri-Rao product and transposed Khatri-Rao product, respectively. $[\mathbf{A}]_{i,j}$ represents the $(i,j)$-th element of matrix $\mathbf{A}$, and $[\mathbf{a}]_{i}$ represents the $i$-th element of vector or set $\mathbf{a}$. Besides, the operator $\text{Diag}(\mathbf{x})$ constructs a diagonal matrix whose diagonal elements are given by the elements of the vector $\mathbf{x}$. The notation $a:b$ denotes the sequence $\{a,a+1,\dots,b\}$ with $a\leq b$. The imaginary unit is denoted as $\jmath=\sqrt{-1}$.
@@ II-A Signal and Channel Model
As depicted in Fig. 2, we consider a RIS-assisted wideband THz MIMO system employing hybrid beamforming architecture that integrates both analog and baseband processing. Assuming that the BS equipped with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}$ (where $N_{\mathrm{RF}}<N_{\mathrm{T}}$) RF chains is deployed to serve $K$ single-antenna UEs. The RIS employs a uniform linear array (ULA) comprised of $N_{\mathrm{R}}$ reflecting elements, which is manipulated by a smart controller operated by the BS. Besides, the orthogonal frequency division multiplexing (OFDM) system is considered with $M$ SCs, central frequency $f_{c}$ and bandwidth $B$.
Figure 2: Illustration of RIS-assisted THz communication system with beam split effect.
@@
In this paper, we assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [12]. For each coherence block, channel estimation is first performed during $T$ subframes, and downlink data transmission is then carried out during the remaining $D-T$ subframes. Moreover, each subframe of the downlink channel estimation period is divided into $P$ time slots, with $P$ baseband beamformers operating alternatively across different time slots. Additionally, we assume that the direct BS-UE links are blocked by obstacles [20] [^1] The proposed scheme can be easily extended to one scenario where direct BS-UE and cascaded BS-RIS-UE links coexist. The direct channel can be estimated via conventional channel estimation approaches by turning off RIS.. Therefore, we mainly focus on the downlink channel estimation of the cascaded BS-RIS-UE channel.
@@
As a cost-performance trade-off, hybrid beamforming architecture is typically employed in THz MIMO systems[15]. During $P$ successive time slots, the frequency-dependent baseband beamformer vectors $\left\{\mathbf{w}_{t,p}^{\mathrm{BB}}[m]\right\}_{p=1}^{P}\in\mathbb{C}^{N_{\mathrm{RF}}}$ are sequentially employed combining with the frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$. Each entry of $\mathbf{W}^{\mathrm{RF}}$ is subject to the unit-modulus constraint, i.e., $|[\mathbf{W}^{\mathrm{RF}}]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ for $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$. We assume that the analog beamformer remains unchanged over multiple channel coherence blocks due to the relatively static nature of the RIS physical location compared to the dynamic channel coefficients [12] [^2] Note that the angular information remains unknown for channel estimation.. Therefore, we define $\mathbf{w}_{t,p}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{w}_{t,p}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ as the hybrid beamformer vector. Let $\mathbf{G}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ and $\mathbf{h}_{k}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}}$ denote the channel from BS to RIS and RIS to $k$-th UE at the $m$-th SC, respectively. At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by $k$-th UE at the $m$-th SC can be expressed as
$\ y_{k,t,p}[m]=\!\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t}[m]\!+\!n_{k,t,p}[m],$ (1)
@@
where $n_{k,t,p}[m]\sim\mathcal{CN}\left(0,2\varepsilon\right)$ is the complex additive white Gaussian noise (AWGN). Besides, $\boldsymbol{\theta}_{t}=\left[e^{\jmath\theta_{t,1}},\cdots,e^{\jmath\theta_{t,N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS at subframe $t$ with a constant reflection amplitude constraint [27].
@@
Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
@@
$\mathbf{G}[m]=\sum_{l=1}^{L}{\alpha_{l}\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{l})\mathbf{a}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\varphi}_{l}\right)e^{-\jmath 2\pi\tau_{l}f_{m}}},$ (4)
where $L$ represents the number of propagation paths between BS and RIS, $\alpha_{l}$, $\tau_{l}$, $\tilde{\psi}_{l}$ and $\tilde{\varphi}_{l}$ represent the complex gain, the time delay, the physical DoA at BS and DoD at RIS for the $l$-th path, respectively. Besides, $f_{m}$ denotes the frequency of the $m$-th SC, which is given by
@@
$\mathbf{a}_{\mathrm{T},m}(\tilde{u})\triangleq\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-\jmath 2\pi\sin\left(\tilde{u}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
$\mathbf{a}_{\mathrm{R},m}(\tilde{v})\triangleq\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-\jmath 2\pi\sin\left(\tilde{v}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th SC. $d_{\mathrm{BS}}$ and $d^{\mathrm{RIS}}$ denote the antenna spacing of BS and RIS, respectively. Besides, $\mathbf{n}_{t}=[0,1,\dots,N_{\mathrm{T}}-1]^{\mathsf{T}}$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]^{\mathsf{T}}$, respectively. By defining the angle vectors $\tilde{\boldsymbol{\psi}}=[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be rewritten as